Robin is having her yard landscaped. She obtained an estimate from two landscaping companies. Company A gave an estimate of $200 for material and equipmwnt rental plus $50 per hour for labor. Company B gave her an estimate of $250 for material and equipment rental plus $40 per hour for labor. Determine how many hours of labor will be required for the two companies to cost the same.
Answer by nerdybill(7384)

Robin is having her yard landscaped. She obtained an estimate from two landscaping companies. Company A gave an estimate of $200 for material and equipmwnt rental plus $50 per hour for labor. Company B gave her an estimate of $250 for material and equipment rental plus $40 per hour for labor. Determine how many hours of labor will be required for the two companies to cost the same.
.
Let h = hours of labor
Company A cost: 200+50h
Company B cost: 250+40h
.
Set them equal and solve for h:
200+50h = 250+40h
50h = 50+40h
10h = 50
h = 5 hours